(A) What is the rms speed of nitrogen molecules contained in a 8.9 m3 volume at 3.1 atm if the total amount of nitrogen is 2300 mol ?

(B) How many atoms are there in a 4.5-g copper coin?
Answer #1

a)

temperature

T = PV/(nR) = 3.1* 1.01* 10^5 *8.9 / (2300* 8.314)

T = 144.28 K

now expression for velocity

v^2 = 3 RT / m = 3* 8.314* 144.28 / 0.014

v = 507 m/s

====

b)

No of atoms = given mass / molar mass * Av No

N = 4.5 / 63* 6.022* 10^23 =4.3* 10^22

====

Comment before rate in case any doubt, will reply for sure.. goodluck